OpenStudy (anonymous):

What would the unit rates be for a temperature change of 20 degrees in 4 hours?

OpenStudy (anonymous):

Well divide 20 by 4 to get 5 degrees an hour

OpenStudy (igreen):

Divide to find out how the degrees change in 1 hour..

OpenStudy (anonymous):

This one is a bit tricky since it shows (blank box) degrees/hour (blank box) hours/degree

OpenStudy (igreen):

Degrees/hour(THIS GOES IN THE FIRST BLANK BOX) = 20/4 = ? hours/degree(THIS GOES IN THE SECOND BLANK BOX) = 4/20 = ?

OpenStudy (anonymous):

Thanks soo much @iGreen !! WIN MEDAL!!

OpenStudy (anonymous):

5 and 0.25?

OpenStudy (igreen):

4 / 20 = 0.20 not 0.25 20 / 4 = 5

OpenStudy (anonymous):

OOps :(